OpenStudy (anonymous):

What is 3 16/25 as a percent?

OpenStudy (anonymous):

3*25= =75 75+16=91 91 is your answer (This is probably the way your teachers want it : 91/100=.91 .91*100= 91% so fraction :3 16/25 decimal .91 percentage 91%)

OpenStudy (anonymous):

Umm... Thats not even one of the choices. The choices are 3.64% 3.80% 364% 380%

OpenStudy (anonymous):

Then its 3.64% . XD I forgot to divide 91 by 25 . sorry !